QUESTION 1

  1. Tom, age 13, is claimed as a dependent by his parents. Tom has unearned income of $3,400 and $300 of income from mowing lawns in the neighborhood. If the first $2,600 of Tom's net unearned income is taxed at 10%, what is Tom’s 2019 income tax liability?

a.

None of these

b.

$260

c.

$370

d.

$0

e.

$459

QUESTION 2

  1. The 3.8 percent ACA Medicare tax does not apply to:

a.

Wages

b.

Interest

c.

Capital gains

d.

Dividends

QUESTION 3

  1. Which one of the following entities cannot use the cash method for tax purposes in 2019?

a.

A small sole proprietorship with $150,000 in average annual gross receipts.

b.

A law firm operating as a personal service corporation with $5.4 million in average annual gross receipts.

c.

All of these may use the cash method.

d.

A corporation with $24 million in average annual gross receipts and 200 employees.

e.

A large almond farm with $40 million in average annual gross receipts.

  1. Assume Karen is 12 years old and her only income is $5,500 of interest income from a bank account with money her parents have given her to save for college. If Karen does not elect to include her income on her parents’ return, what is her income tax liability?

a.

$538

b.

$963

c.

$550

d.

$445

QUESTION 5

  1. William, a cash-basis sole proprietor, had the following receipts and disbursements for the current year:

Gross receipts

$50,000

Cost of sales

30,000

Other operating expenses

5,000

Personal medical expenses

600

For the current year, what amount should William report as net earnings from self-employment?

a.

$15,000

b.

$14,000

c.

None of these

d.

$20,000

e.

$13,400

QUESTION 6

  1. In cash basis accounting, for tax purposes:

a.

Income is generally recognized when it is actually or constructively received and expenses are generally recognized when they are paid.

b.

Income is recognized when it is actually or constructively received and expenses are recognized when they are actually or constructively incurred, regardless of when paid.

c.

Income is recognized when it is earned regardless of when received and expenses are recognized when they are actually or constructively incurred.

d.

The cash basis is not allowed for businesses reported on Schedule C.

QUESTION 7

  1. Which of the following is not an acceptable method of accounting under the tax law?

a.

All of these are acceptable

b.

The hybrid method

c.

The accrual method

d.

The cash method

e.

None of these are acceptable

QUESTION 8

  1. A parent may elect to include a child’s income in the parent’s return if:

a.

All of these must be met for a parent to elect to include a child’s income in the parent’s return.

b.

The child’s income is only from interest and dividend distributions.

c.

No estimated tax has been paid in the name of the child and the child is not subject to backup withholding.

d.

The child’s gross income is more than $1,100 and less than $11,000.

QUESTION 9

  1. Becky is a cash basis taxpayer with the following transactions during her calendar tax year:

Cash basis revenue

$54,000

Cash basis expenses, except rent

25,000

Rent expense (paid on December 1) for use of a building for 24 months

24,000

What is the amount of Becky's taxable income from her business for this tax year?

a.

None of these

b.

$28,000

c.

$27,500

d.

$11,000

e.

$7,000 loss

QUESTION 10

  1. The Dot Corporation has changed its year-end from a calendar year-end to August 31. The income for its short period from January 1 to August 31 is $54,000. The tax for this short period is:

a.

$6,250

b.

$2,040

c.

$8,667

d.

$11,340
Answer #1

Dear Student,

As per the HOMEWORKLIB POLICY, only the first four questions should be answered. Kindly take note of it.

Part 1

Answer is option B

B. $260

Taxable income = unearned income + earned income – standard deduction = 3400+300-1100 =2600

Tax liability = 2600*10% = $260

Part 2

Answer is option A

A. Wages

The 3.8 percent ACA Medicare tax does not apply to earned income. It applies only to net investment income

Part 3

Answer is option C

C. All of these may use the cash method.

All given entities meets the requirements and are eligible to use cash method for tax purpose in 2019

Part 4

Answer is option A

A. $538

Income tax liability = 260+(((5500-2200)-2600)*24%)+(1100*10%)